OpenStudy (anonymous):

HELP FAST I WILL FAN AND MEDAL How many different lunch combinations can be made from two sandwich's choices, and three beverage choices if you choose one sandwich's , one side and one beverage A. 8 B. 9 C. 18 D. 11

OpenStudy (anonymous):

@agent0smith @BSwan @ganeshie8 @hippo4 @iambatman @ikatouni @iPwnBunnies @JayBaybie @mlb004 @SolomonZelman @Whitemonsterbunny17 @Hero @Zarkon @Destinymasha @myininaya

whitemonsterbunny17 (whitemonsterbunny17):

How many side choices are there?

OpenStudy (anonymous):

i think 2

whitemonsterbunny17 (whitemonsterbunny17):

For this, you would have to basically draw it out and count the combinations... Something like this: S1 (Sandwich 1), S2 (Sandwich 2), D1 (Drink 1), D2 (Drink 2), D3 (Drink 3), Si1 (Side 1), Si2 (Side 2) Combination 1: S1 D1 Si1 Combination 2: S1 D2 Si1 Combination 3: S1 D3 Si1 Combination 4: S1 D1 Si2 Combination 5: S1 D2 Si2 Combination 6: S1 D3 Si2 Combination 7: S2 D1 Si1 Combination 8: S2 D2 Si1 Combination 9: S2 D3 Si1 Combination 10: S2 D1 Si2 Combination 11: S2 D2 Si2 Combination 12: S2 D3 Si2

OpenStudy (anonymous):

hello

whitemonsterbunny17 (whitemonsterbunny17):

Does that make sense?

OpenStudy (anonymous):

yea but what does it mean

whitemonsterbunny17 (whitemonsterbunny17):

You have to write out all of the possible combinations, then count them. As I just did.

whitemonsterbunny17 (whitemonsterbunny17):

I got 12.. but that doesn't match your choices.

whitemonsterbunny17 (whitemonsterbunny17):

Unless the number of sides isn't 2.. then that really changes things.

OpenStudy (anonymous):

oh i dont know what the number of sides are i was looking at the question to find that

OpenStudy (xguardians):

Where is number of sides?

OpenStudy (anonymous):

umm how do you find that

OpenStudy (xguardians):

It's not in the question?

OpenStudy (xguardians):

Actually I can Guarantee you that the answer will be C. without knowing the # of sides.

OpenStudy (anonymous):

i think the side is 1

OpenStudy (xguardians):

2 (# Sandwiches) * 3 (# of Beverages) * 3 (# of sides) = 18. Is the only feasible option. 2 (# Sandwiches) * 3 (# of Beverages) * 1(# of sides) = 6 Not an Option. 2 (# Sandwiches) * 3 (# of Beverages) * 2 (# of sides) = 12. Not an Option Therefore the number of sides is probably three, thus making the answer C-18

OpenStudy (anonymous):

thanxs @Whitemonsterbunny17 and @xGuardians

whitemonsterbunny17 (whitemonsterbunny17):

You're welcome!! *^.^*

OpenStudy (xguardians):

No problem :)